Evaluate JF.dr from (0,0) to (9,3) along the curve y=vx if Ē(x, y, z) = x’yi – xy? j -
Answer #1

(0,0) to 19,3) To yasa F En 1,2) = x2 yi - ny2; 4- 0 = got nagt, y=34 da-got, dy = 300- dy= 222 dx = 2 Joe dy = farzda – , ga